carmelle:

Factor completely 3x^2 − x − 4. (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1)

snowflake0531:

if the 3, in ax^2, was not there... do you know how to factor these kinds of stuff

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 if the 3, in ax^2, was not there... do you know how to factor these kinds of stuff \(\color{#0cbb34}{\text{End of Quote}}\) no.. 😐

snowflake0531:

:/, you're teacher first gave you harder thing before the fundamental stuffO-O oof

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 :/, you're teacher first gave you harder thing before the fundamental stuffO-O oof \(\color{#0cbb34}{\text{End of Quote}}\) wait, actually I think ik how to do it-

snowflake0531:

well, would you know how to simplify each of the choices

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 well, would you know how to simplify each of the choices \(\color{#0cbb34}{\text{End of Quote}}\) idk, im confused

snowflake0531:

do you know how to simplify each of these (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1)

carmelle:

yea, I think so.. correct me if im wrong, is the first one 3x^2 + 11x - 4

snowflake0531:

yes and AZ will take over he's better at explaining this kind of stuff

carmelle:

ok

AZ:

so yeah you could do that method too but let's learn how to factor it the right way in case there isn't any answer choices

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so yeah you could do that method too but let's learn how to factor it the right way in case there isn't any answer choices \(\color{#0cbb34}{\text{End of Quote}}\) ok

AZ:

when we have ax^2 + bx + c where a, b, and c are just the coefficients then when we're factoring we have to find a number that multiplies to a*c and adds up to b

AZ:

so for your question 3x^2 − x − 4 3*(-4) = -12 so what two numbers multiply to give us -12 and add up to give us -1

carmelle:

I-

carmelle:

-6 x 2 = -12

AZ:

forget about the -12 what are all the factors of 12 what two numbers can multiply to 12

carmelle:

im confused :|

AZ:

so what are the factors of 12 like 1 * 12 = 12 2 * 6 = 12 3 * 4 = 12

carmelle:

ohhhh

AZ:

now remember we said \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so for your question 3x^2 − x − 4 3*(-4) = -12 so what two numbers multiply to give us -12 and add up to give us -1 \(\color{#0cbb34}{\text{End of Quote}}\)

AZ:

we need to make one of the numbers negative so that way when we multiply it, we'll get negative 12 and when we add it, we'll get negative 1

AZ:

so do we use -3 and 4 or -4 and 3

carmelle:

-4 and 3

AZ:

good! so now that we have those two numbers we can write 3x^2 - x - 4 that -x we can re-write it as 3x - 4x because that's equal to -x so we have 3x^2 + 3x - 4x - 4 do you know how to factor by grouping?

carmelle:

im not sure

AZ:

so to do it we factor it as halfs

AZ:

so we have 3x^2 + 3x - 4x - 4 so let's factor 3x^2 + 3x and factor -4x -4

carmelle:

3x(x + 1)

AZ:

Good! What about -4x - 4

carmelle:

and -4(x+1)

carmelle:

I have to go :(

AZ:

we're almost done!

carmelle:

ok

AZ:

but we can finish later if you have to go go

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ but we can finish later if you have to go go \(\color{#0cbb34}{\text{End of Quote}}\) lets finish it bc I won't be back any time soon ;-;

AZ:

so we went from 3x^2 + 3x - 4x - 4 to 3x(x + 1) -4(x+1) so now we have 3x(x + 1) -4(x+1) so now think of the (x+1) as an apple or let's use a new letter such as `c` just so that we can see everything more clearly so we have 3x*c -4c can you see it now?

AZ:

can you factor out the c from 3xc - 4c

carmelle:

ah, c(3x-4)

AZ:

exactly!! and we were just using `c` in place of (x+1) since that was so big so essentially c(3x-4) is (x+1)(3x-4)

AZ:

ta-da!

carmelle:

:o tysmm

carmelle:

I get it now-

AZ:

yay!!

AZ:

you're most welcome :D

carmelle:

I feel so dumb 😅

snowflake0531:

everyone does while learning~

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 everyone does while learning~ \(\color{#0cbb34}{\text{End of Quote}}\) yeah B)

AZ:

no no no I was once like you too learning this and now I'm much better at it (Like I even saw the question and was hesitant to take a stab at it because of the 3x^2 so I told snowflake to do it until I remembered and took over so don't feel bad at all!)

carmelle:

xD thx for the encouragement guys :)

carmelle:

or girls 0-0

AZ:

you're welcome :D

snowflake0531:

lmao

carmelle:

We meet again o;

snowflake0531:

xd